OpenStudy (anonymous):

solve for x: 1) ln (3x-5) = ln2 + ln (x-1) 2) 3 log x = log 64 3) log x + log (2x+1) = 1

OpenStudy (anonymous):

1. Get x on one side, use the quotient rule of combining logarithms, and take each side of the equation as a power of "e" 2. Divide both sides by three, and take each side of the equation to the 10th power 3. Use the product rule of logarithms to combine the first two terms on the left, take each side of the equation to the tenth power.

OpenStudy (anonymous):

for the first one. ln(3x-5)=ln2+ln(x-1) you will have the formula that says lna+lnb=ln(ab) so you have ln(3x-5)=ln(2(x-1))---->ln(3x-5)=ln(2x-2)----->3x-5=2x-2--->x=3